OpenStudy (anonymous):

Let f ( x) = 4 - x and g ( x) = x + 4. (f/g)(x)=

OpenStudy (akshay_budhkar):

(f/g)(x)=f(x)/g(x) therefore your answer is \[(f/g)x={(4-x)\over(x+4)}\]

OpenStudy (anonymous):

Thanks so much Akshay :)

OpenStudy (akshay_budhkar):

your welcome :)